# Byzantine-resilient Federated Learning With Adaptivity to Data Heterogeneity

### v3

@@ IV Analysis of Convergence and Robustness
In this section, we provide theoretical analysis of the robustness and convergence of the RAGA under Byzantine attacks under two cases of loss function’s convexity.
IV-A Case \@slowromancapi@: With Lipschitz Continuity Only for Loss Function
We first inspect the case with Assumptions 1, 3, 4 and 5 imposed, which allows the loss function to be non-convex. By defining
@@ IV-A Case \@slowromancapi@: With Lipschitz Continuity Only for Loss Function
$p^{t}=\left\{\begin{aligned} &{0,}&&{\eta^{t}\leqslant\frac{1}{L}}\\ &{1,}&&{\eta^{t}>\frac{1}{L}}\end{aligned}\right.$ (18)
@@
$\displaystyle\quad\sum_{t=1}^{T}\frac{\eta^{t}}{\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}\leqslant\frac{2\mathbb{E}\left\{F(\bm{w}^{1})-F(\bm{w}^{T})\right\}}{\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}+\frac{1}{\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}\sum_{t=1}^{T}\eta^{t}\Delta^{t}$
$\displaystyle\quad+\sum_{t=1}^{T}\frac{2(\eta^{t}+p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})\epsilon^{2}}{(2C_{\alpha}-1)^{2}\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}+\sum_{t=1}^{T}\frac{8(C_{\alpha})^{2}(\eta^{t}\sigma^{2}+(p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})(G^{2}+\sigma^{2}))}{(2C_{\alpha}-1)^{2}\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}$ (19)
@@
$\displaystyle\Delta^{t}=$ $\displaystyle\frac{8C_{\alpha}}{(2C_{\alpha}-1)^{2}}\sum_{m\in\mathcal{M}\setminus\mathcal{B}}\Bigg{(}\frac{2\alpha_{m}L^{2}(G^{2}+\sigma^{2})}{K^{t}}\sum_{k=2}^{K^{t}}(k-1)\cdot\sum_{i=1}^{k-1}(\eta_{m}^{t,i})^{2}+\frac{2\alpha_{m}}{K^{t}}\sum_{k=1}^{K^{t}}\theta_{m}^{t,k-1}\Bigg{)},$ (20)
and
$\displaystyle\theta_{m}^{t,k}\triangleq\left\lVert\nabla F_{m}(\bm{w}_{m}^{t,k})-\nabla F(\bm{w}_{m}^{t,k})\right\rVert^{2},$ (21)
which actually represents the data heterogeneity of $k$th update in $t$ iteration for user $m$.
@@
Please refer to Appendix A. ∎
@@
    $\lim_{T\rightarrow\infty}\sum_{t=1}^{T}\eta^{t}\sum_{m\in\mathcal{M}\setminus\mathcal{B}}\sum_{i=1}^{K^{t}}(\eta_{m}^{t,i})^{2}<\infty$, and there are only finitely many $t$ such that $p^{t}=1$, together with the prerequisite condition of Theorem 1, i.e., $C_{\alpha}>0.5$, it can be derived that the term $\sum_{t=1}^{T}\frac{\eta^{t}}{\sum_{t^{\prime}=1}^{T}\eta^{t}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$ will converge to $\mathcal{O}\left(\max\{\theta_{m}^{t,k}\}+\epsilon^{2}+\sigma^{2}\right)$.
  • To be concrete, we set $\eta_{m}^{t,k}=\eta^{t}\leqslant\frac{1}{L}$, $K^{t}=K$, and $\theta_{m}^{t,k}\leqslant\theta$. Then inequality (1) dwells into
    $\displaystyle\quad\sum_{t=1}^{T}\frac{\eta^{t}}{\sum_{t^{\prime}=1}^{T}\eta^{t}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}\leqslant\frac{8(C_{\alpha})^{2}L^{2}(G^{2}+\sigma^{2})(K-1)(2K-1)\sum_{t=1}^{T}(\eta^{t})^{3}}{3(2C_{\alpha}-1)^{2}\sum_{t=1}^{T}\eta^{t}}$
    $\displaystyle+\frac{2\mathbb{E}\left\{F(\bm{w}^{1})-F(\bm{w}^{T})\right\}}{\sum_{t=1}^{T}\eta^{t}}+\frac{16(C_{\alpha})^{2}\theta+2\epsilon^{2}+8(C_{\alpha})^{2}\sigma^{2}}{(2C_{\alpha}-1)^{2}}$ (22)
@@
    $\frac{16(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}\theta+\frac{2}{(2C_{\alpha}-1)^{2}}\epsilon^{2}+\frac{8(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}\sigma^{2}$ (23)
@@
  • To further concrete convergence rate, we set $\eta^{t}=\frac{1}{L}\frac{1}{t^{1/3+\delta}}$, where $\delta$ is an arbitrary value lying between $(0,2/3)$, then the right-hand side of (1) is upper bounded by
    $\displaystyle\mathcal{O}\left(\frac{8(C_{\alpha})^{2}(G^{2}+\sigma^{2})(K-1)(2K-1)(1+3\delta-3\delta T^{-3\delta})}{3(2C_{\alpha}-1)^{2}T^{2/3-\delta}}\right)$
    $\displaystyle+\mathcal{O}\left(\frac{2L\mathbb{E}\left\{F(\bm{w}^{1})-F(\bm{w}^{T})\right\}}{T^{2/3-\delta}}\right)+\frac{16(C_{\alpha})^{2}\theta+2\epsilon^{2}+8(C_{\alpha})^{2}\sigma^{2}}{(2C_{\alpha}-1)^{2}}$ (24)
@@
IV-B Case \@slowromancapii@: With Lipschitz Continuity and Strong Convexity for Loss Function
For the case with Assumptions 1, 2, 3, 4 and 5 imposed, which requires the loss function to be not only Lipschitz but also strongly-convex, the following results can be anticipated.
@@ IV-B Case \@slowromancapii@: With Lipschitz Continuity and Strong Convexity for Loss Function
With $1\displaystyle 0<\lambda^{t}<$ 1 and $C_{\alpha}>0.5$, the optimality gap $\mathbb{E}\left\{F(\bm{w}^{T})-F(\bm{w}^{*})\right\}$ is upper bounded in (2) as follows,
@@
$\displaystyle\leqslant\frac{L}{2}\mathbb{E}\left\{\left\lVert\bm{w}^{1}-\bm{w}^{*}\right\rVert^{2}\right\}\prod_{t=1}^{T-1}\gamma^{t}+\frac{L}{2}\sum_{t=1}^{T-1}\frac{(\eta^{t})^{2}}{\lambda^{t}}\Bigg{(}\Delta^{t}+\frac{8\sigma^{2}(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}\Bigg{)}\cdot\prod_{i=t}^{T-1}(\gamma^{i+1})^{q^{i+1}}$ (25)
@@
$q^{t}=\left\{\begin{aligned} &{0,}&&{t=T}\\ &{1,}&&{t<T}\end{aligned}\right.$
@@
Please refer to Appendix B. ∎
@@
  • For general case, as long as $\eta_{m}^{t,k}\propto T^{-\beta},\beta>0$, $\lambda^{t}=\rho^{t}\eta^{t}$ with $\rho^{t}<2\mu$ and $\eta^{t}<\frac{2\mu-\rho^{t}}{L^{2}}$, $1\displaystyle 0<\gamma^{t}<$ 1, and the preconditions listed in Theorem 2 hold, the convergence of our proposed RAGA can be ensured and the optimality gap is at the scale $\mathcal{O}\left(\max\{\theta_{m}^{t,k}\}+\epsilon^{2}+\sigma^{2}\right)$.
  • To concrete the optimality gap, we suppose $\eta_{m}^{t,k}=\frac{1}{T}\eta$, $\eta^{t}=\eta$, $\gamma^{t}=\gamma$, $K^{t}=K$, $\lambda^{t}=\mu\eta^{t}=\mu\eta$, $\eta^{t}<\frac{\mu}{L^{2}}$, and $\theta_{m}^{t,k}\leqslant\theta$, then there is
    $\displaystyle\quad\mathbb{E}\left\{F(\bm{w}^{T})-F(\bm{w}^{*})\right\}\leqslant\frac{L(\gamma)^{T-1}}{2}\mathbb{E}\left\{\left\lVert\bm{w}^{1}-\bm{w}^{*}\right\rVert^{2}\right\}$
    $\displaystyle+\frac{4L^{3}(K-1)(2K-1)(C_{\alpha})^{2}(G^{2}+\sigma^{2})}{3\mu(2C_{\alpha}-1)^{2}T^{2}}\frac{1-(\gamma)^{T}}{1-\gamma}(\eta)^{3}+\frac{8L(C_{\alpha})^{2}\theta+\epsilon^{2}L+4L(C_{\alpha})^{2}\sigma^{2}}{\mu(2C_{\alpha}-1)^{2}}\frac{1-(\gamma)^{T}}{1-\gamma}\eta.$ (26)
@@
    $\frac{1-\mu\eta}{\mu-L^{2}\eta}\left(\frac{8L(C_{\alpha})^{2}}{\mu(2C_{\alpha}-1)^{2}}\theta+\frac{L}{\mu(2C_{\alpha}-1)^{2}}\epsilon^{2}+\frac{4L(C_{\alpha})^{2}}{\mu(2C_{\alpha}-1)^{2}}\sigma^{2}\right)$ (27)
@@
    $\displaystyle+\left(\frac{L(\gamma)^{T-1}}{2}\mathbb{E}\left\{\left\lVert\bm{w}^{1}-\bm{w}^{*}\right\rVert^{2}\right\}\right)+\frac{8L(C_{\alpha})^{2}\theta+\epsilon^{2}L+4L(C_{\alpha})^{2}\sigma^{2}}{\mu(2C_{\alpha}-1)^{2}}\frac{1-(\gamma)^{T}}{1-\gamma}\eta$ (28)

### v4

@@ IV Analysis of Convergence and Robustness
In this section, we provide theoretical analysis of the robustness and convergence of the RAGA under Byzantine attacks under two cases of loss function’s convexity. Below, we first present the necessary assumptions.
IV-A Assumption
For all honest users $m\in\mathcal{M}\setminus\mathcal{B}$, we have the following assumptions.
Assumption 1 (Lipschitz Continuity)
$\displaystyle f(\bm{w}_{1},\bm{s})-f(\bm{w}_{2},\bm{s})$ $\displaystyle\leqslant\left\langle\nabla f(\bm{w}_{2},\bm{s}),\bm{w}_{1}-\bm{w}_{2}\right\rangle$
$\displaystyle\qquad\qquad\quad+\frac{L}{2}\left\lVert\bm{w}_{1}-\bm{w}_{2}\right\rVert^{2},$ (10)
which is equivalent to the following inequality,
With (7) or (31), it can be derived that the global loss function $F(\bm{w})$ and local loss function $F_{m}(\bm{w})$ both have $L$-Lipschitz continuity.
Assumption 2 (Strongly-Convex)
Similarly, with (32), $F(\bm{w})$ and $F_{m}(\bm{w})$ can be derived to be both $\mu$ strongly-convex.
Assumption 3 (Local Unbiased Gradient)
Suppose a subset of dataset $\mathcal{S}_{m}$ is selected randomly, which is denote as $\xi_{m}$, define
the unbiased gradient [37, 38] implies that
$\mathbb{E}\left\{\nabla F_{m}(\bm{w};\xi_{m})\right\}=\nabla F_{m}(\bm{w}).$ (14)
Assumption 4 (Bounded Inner Variance)
For each honest user $m$ and $\bm{w}\in\mathbb{R}^{p}$, the variance of stochastic local gradient $\nabla F_{m}(\bm{w};\xi_{m})$ is upper-bounded [37, 38] by
$\displaystyle\text{Var}(\nabla F_{m}(\bm{w};\xi_{m}))\leqslant\sigma^{2}.$ (15)
Assumption 5 (Data Heterogeneity)
For each honest user $m$ and $\bm{w}\in\mathbb{R}^{p}$, the data heterogeneity can be defined as
$\displaystyle\theta_{m}=\left\lVert\nabla F_{m}(\bm{w})-\nabla F(\bm{w})\right\rVert.$ (16)
We also assume the data heterogeneity is bounded [37, 38], which implies
$\displaystyle\theta_{m}\leqslant\theta.$ (17)
Assumption 6 (Bounded Gradient)
For each honest user $m$ and $\bm{w}\in\mathbb{R}^{p}$, the ideal local gradient $\nabla F_{m}(\bm{w})$ is upper-bounded [37, 38] by
$\displaystyle\left\lVert\nabla F_{m}(\bm{w})\right\rVert\leqslant G.$ (18)
IV-B Convergence Analysis
In this subsection, we present the analytical results of convergence on our proposed RAGA. All proofs are deferred to Appendices B and C.
IV-B1 Case I: With Lipschitz Continuity Only for Loss Function
We first inspect the case with Assumptions 7, 9, 10, 11 and 12 imposed, which allows the loss function to be non-convex. By defining
@@ IV-B1 Case I: With Lipschitz Continuity Only for Loss Function
$p^{t}=\left\{\begin{aligned} &{0,}&&{\eta^{t}\leqslant\frac{1}{L}}\\ &{1,}&&{\eta^{t}>\frac{1}{L}}\end{aligned}\right$., (20)
@@
$\displaystyle\quad\sum_{t=1}^{T}\frac{\eta^{t}}{\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$
$\displaystyle\leqslant\frac{2\mathbb{E}\left\{F(\bm{w}^{1})-F(\bm{w}^{T})\right\}}{\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}+\frac{1}{\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}\sum_{t=1}^{T}\eta^{t}\Delta^{t}$
$\displaystyle\quad+\sum_{t=1}^{T}\frac{2(\eta^{t}+p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})\epsilon^{2}}{(2C_{\alpha}-1)^{2}\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}}$
$\displaystyle\quad+\sum_{t=1}^{T}\frac{8(C_{\alpha})^{2}(\eta^{t}\sigma^{2}+(p^{t}(\eta^{t})^{2}L-p^{t}\eta^{t})(G^{2}+\sigma^{2}))}{(2C_{\alpha}-1)^{2}\sum_{{t^{\prime}}=1}^{T}\eta^{t^{\prime}}},$ (21)
@@
$\displaystyle\Delta^{t}=$ $\displaystyle\frac{8C_{\alpha}}{(2C_{\alpha}-1)^{2}}\sum_{m\in\mathcal{M}\setminus\mathcal{B}}\Bigg(\frac{2\alpha_{m}L^{2}(G^{2}+\sigma^{2})}{K^{t}}\sum_{k=2}^{K^{t}}(k-1)$
$\displaystyle\quad\cdot\sum_{i=1}^{k-1}(\eta_{m}^{t,i})^{2}+2\alpha_{m}\theta^{2}\Bigg).$ (22)
@@
Please refer to Appendix B. ∎
@@
    $\lim_{T\rightarrow\infty}\sum_{t=1}^{T}\eta^{t}\sum_{m\in\mathcal{M}\setminus\mathcal{B}}\sum_{i=1}^{K^{t}}(\eta_{m}^{t,i})^{2}<\infty$, and there are only finitely many $t$ such that $p^{t}=1$, together with the prerequisite condition of Theorem 3, i.e., $C_{\alpha}>0.5$, it can be derived that the term $\sum_{t=1}^{T}\frac{\eta^{t}}{\sum_{t^{\prime}=1}^{T}\eta^{t}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$ will converge to $\mathcal{O}\left(\theta^{2}+\epsilon^{2}+\sigma^{2}\right)$.
  • To be concrete, we set $\eta_{m}^{t,k}=\eta^{t}\leqslant\frac{1}{L}$ and $K^{t}=K$. Then inequality (3) dwells into
    $\displaystyle\qquad\sum_{t=1}^{T}\frac{\eta^{t}}{\sum_{t^{\prime}=1}^{T}\eta^{t}}\mathbb{E}\left\{\left\lVert\nabla F(\bm{w}^{t})\right\rVert^{2}\right\}$
    $\displaystyle\leqslant\frac{8(C_{\alpha})^{2}L^{2}(G^{2}+\sigma^{2})(K-1)(2K-1)\sum_{t=1}^{T}(\eta^{t})^{3}}{3(2C_{\alpha}-1)^{2}\sum_{t=1}^{T}\eta^{t}}$
    $\displaystyle+\frac{2\mathbb{E}\left\{F(\bm{w}^{1})-F(\bm{w}^{T})\right\}}{\sum_{t=1}^{T}\eta^{t}}+\frac{16(C_{\alpha})^{2}\theta^{2}+2\epsilon^{2}+8(C_{\alpha})^{2}\sigma^{2}}{(2C_{\alpha}-1)^{2}},$ (23)
@@
    $\frac{16(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}\theta^{2}+\frac{2}{(2C_{\alpha}-1)^{2}}\epsilon^{2}+\frac{8(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}\sigma^{2},$ (24)
@@
  • To further concrete convergence rate, we set $\eta^{t}=\frac{\eta_{0}}{L}\frac{1}{(t+c)^{1/3+\delta}}$, where $\delta$ is an arbitrary value lying between $(0,2/3)$, $\eta_{0}<1$ and $c>0$, the right shift introduced by adding $c>0$ to the denominator is intended to mitigate fluctuations during the initial training iterations, then the right-hand side of (21) of the revised manuscript is upper bounded by
    $\displaystyle\mathcal{O}\left(\frac{8\eta_{0}^{2}(C_{\alpha})^{2}(G^{2}+\sigma^{2})(K-1)(2K-1)(1+3\delta-3\delta T^{-3\delta})}{3(2C_{\alpha}-1)^{2}T^{2/3-\delta}}\right)$
    $\displaystyle+\mathcal{O}\left(\frac{2L\mathbb{E}\left\{F(\bm{w}^{1})-F(\bm{w}^{T})\right\}}{\eta_{0}T^{2/3-\delta}}\right)$
    $\displaystyle+\frac{16(C_{\alpha})^{2}\theta^{2}+2\epsilon^{2}+8(C_{\alpha})^{2}\sigma^{2}}{(2C_{\alpha}-1)^{2}},$ (25)
@@
IV-B2 Case II: With Lipschitz Continuity and Strong Convexity for Loss Function
For the case with Assumptions 7, 8, 9, 10, 11 and 12 imposed, which requires the loss function to be not only Lipschitz but also strongly-convex, the following results can be anticipated.
@@ IV-B2 Case II: With Lipschitz Continuity and Strong Convexity for Loss Function
With $0<\lambda^{t}<1$ and $C_{\alpha}>0.5$, the optimality gap $\mathbb{E}\left\{F(\bm{w}^{T})-F(\bm{w}^{*})\right\}$ is upper bounded in (4) as follows,
@@
$\displaystyle\leqslant\frac{L}{2}\mathbb{E}\left\{\left\lVert\bm{w}^{1}-\bm{w}^{*}\right\rVert^{2}\right\}\prod_{t=1}^{T-1}\gamma^{t}+\frac{L}{2}\sum_{t=1}^{T-1}\frac{(\eta^{t})^{2}}{\lambda^{t}}\Bigg(\Delta^{t}+$
$\displaystyle\quad\frac{8\sigma^{2}(C_{\alpha})^{2}}{(2C_{\alpha}-1)^{2}}+\frac{2\epsilon^{2}}{\left(2C_{\alpha}-1\right)^{2}}\Bigg)\cdot\prod_{i=t}^{T-1}(\gamma^{i+1})^{q^{i+1}},$ (26)
@@
$q^{t}=\left\{\begin{aligned} &{0,}&&{t=T}\\ &{1,}&&{t<T}\end{aligned}\right$.,
@@
Please refer to Appendix C. ∎
@@
  • For general case, as long as $\eta_{m}^{t,k}\propto T^{-\beta},\beta>0$, $\lambda^{t}=\rho^{t}\eta^{t}$ with $\rho^{t}<2\mu$ and $\eta^{t}<\frac{2\mu-\rho^{t}}{L^{2}}$, $0<\gamma^{t}<1$, and the preconditions listed in Theorem 4 hold, the convergence of our proposed RAGA can be ensured and the optimality gap is at the scale $\mathcal{O}\left(\theta^{2}+\epsilon^{2}+\sigma^{2}\right)$.
  • To concrete the optimality gap, we suppose $\eta_{m}^{t,k}=\frac{1}{T}\eta$, $\eta^{t}=\eta$, $\gamma^{t}=\gamma$, $K^{t}=K$, $\lambda^{t}=\mu\eta^{t}=\mu\eta$, and $\eta^{t}<\frac{\mu}{L^{2}}$, then there is
    $\displaystyle\quad\mathbb{E}\left\{F(\bm{w}^{T})-F(\bm{w}^{*})\right\}$
    $\displaystyle\leqslant\frac{L(\gamma)^{T-1}}{2}\mathbb{E}\left\{\left\lVert\bm{w}^{1}-\bm{w}^{*}\right\rVert^{2}\right\}$
    $\displaystyle+\frac{4L^{3}(K-1)(2K-1)(C_{\alpha})^{2}(G^{2}+\sigma^{2})}{3\mu(2C_{\alpha}-1)^{2}T^{2}}\frac{1-(\gamma)^{T}}{1-\gamma}(\eta)^{3}$
    $\displaystyle+\frac{8L(C_{\alpha})^{2}\theta^{2}+\epsilon^{2}L+4L(C_{\alpha})^{2}\sigma^{2}}{\mu(2C_{\alpha}-1)^{2}}\frac{1-(\gamma)^{T}}{1-\gamma}\eta.$ (27)
@@
    $\displaystyle\frac{1-\mu\eta}{\mu-L^{2}\eta}\left(\frac{8L(C_{\alpha})^{2}}{\mu(2C_{\alpha}-1)^{2}}\theta^{2}+\frac{L}{\mu(2C_{\alpha}-1)^{2}}\epsilon^{2}\right$.
    $\displaystyle\left.+\frac{4L(C_{\alpha})^{2}}{\mu(2C_{\alpha}-1)^{2}}\sigma^{2}\right),$ (28)
@@
    $\displaystyle+\left(\frac{L(\gamma)^{T-1}}{2}\mathbb{E}\left\{\left\lVert\bm{w}^{1}-\bm{w}^{*}\right\rVert^{2}\right\}\right)$
    $\displaystyle+\frac{8L(C_{\alpha})^{2}\theta^{2}+\epsilon^{2}L+4L(C_{\alpha})^{2}\sigma^{2}}{\mu(2C_{\alpha}-1)^{2}}\frac{1-(\gamma)^{T}}{1-\gamma}\eta,$ (29)